OpenStudy (anonymous):

The quadractic x(squared) +5x-14 has what type of roots?

OpenStudy (rational):

\(\large x^2 + 5x - 14\)

OpenStudy (rational):

find the discriminant

OpenStudy (rational):

\(\large D = \sqrt{b^2 - 4ac}\)

OpenStudy (anonymous):

Yes , this is confusing :(

OpenStudy (rational):

do you know the standard form of quadratic equation ?

OpenStudy (anonymous):

Yes

OpenStudy (rational):

\(\large ax^2 + bx + c\)

OpenStudy (rational):

compare your equation with above, and figure out below : a = ? b = ? c = ?

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

I have the discriminant but how do I figure out the roots ?

OpenStudy (rational):

use below : if \(D > 0\) then, "two different real roots" if \(D < 0 \) then, "two different complex roots" if \(D = 0\) then, "two SAME rational roots"

OpenStudy (rational):

wat did u get for \(D\) ?

OpenStudy (anonymous):

-31

OpenStudy (rational):

nope.. try again

OpenStudy (anonymous):

This is hard :(

OpenStudy (rational):

\(\large x^2 + 5x - 14 \) a = 1 b = 5 c = -14

OpenStudy (rational):

right ?

OpenStudy (anonymous):

Yes

OpenStudy (rational):

\(\large D = \sqrt{b^2 - 4ac}\) \(\large D = \sqrt{(5)^2 - 4(1)(-14)}\) \(\large D = \sqrt{25 + 56}\) \(\large D = \sqrt{81}\) \(\large D = 9\)

OpenStudy (rational):

fine ?

OpenStudy (anonymous):

So the discriminant is 9

OpenStudy (rational):

yes :)

OpenStudy (rational):

so the roots are "two real roots" but they will be "rational" because 81 is a perfect square

OpenStudy (rational):

"two different rational roots"

OpenStudy (rational):

let me correct a mistake ive started wid : \(\large D = b^2 - 4ac\) square-root sign is not there ok ?

OpenStudy (anonymous):

Thank you ! And I just starting learning this today.. I think I should catch on

OpenStudy (rational):

you wlcme :) good luck !!